a cyclist is going from mile mark 20 to mile mark 50 in 30 mins. what is the cyclist speed ? can we know the cyclist acceleration ? why? what is the cyclists speed min to second ?
1
Expert's answer
2012-10-26T10:11:50-0400
The average speed of a cyclist is

S = (50[mile]-20[mile])/30[min] = 1[mile/min] = 1609.344/60[m/s] = 26.8224[m/s].

We can not find the cyclist's acceleration cause we don't have enough information about the changes of his speed.